OpenStudy (anonymous):

1. is y =1/2x^2 Quadratic? Explain your reasoning

OpenStudy (anonymous):

exponent is 2 form of quadratic is y=ax^2 +bx + c

OpenStudy (anonymous):

a=1/2 b and c = 0 so yes

OpenStudy (mathmagician):

but if you mean y=1/(2x^2), then it is hyperbolic

OpenStudy (anonymous):

thanks both of you :)